OpenStudy (anonymous):

prime question: f(x)=2e^(x+2)+e^a f`(x)=

OpenStudy (anonymous):

mistake: e^a=> e^9

OpenStudy (anonymous):

f`(x)=2e^(x+2)

OpenStudy (stamp):

\[f(x)=2e^{x+2}+e^9\]\[f'(x)=2e^{x+2}\]

OpenStudy (anonymous):

how did you get this answer?

OpenStudy (stamp):

\[f(x)=2e^{x+2}+e^9\] \[d(2e^{x+2})/dx=2\ d(e^{x+2})/dx\]

OpenStudy (stamp):

u = x+2, du = 1 dx\[e^{x+2}=e^u=du/dx\ e^u=1e^{x+2}\]

OpenStudy (anonymous):

I entered the answer and it told me that it was wrong

OpenStudy (anonymous):

No i am sorry! f`(0)=

OpenStudy (anonymous):

no x!

OpenStudy (stamp):

So plug in 0

OpenStudy (stamp):

You need to get your stuff together, dodo1

OpenStudy (anonymous):

into x?

OpenStudy (stamp):

f'(x) is a function. So yes, f'(0) would be f'(x) with 0 for x.

OpenStudy (anonymous):

so is 2?

OpenStudy (anonymous):

no its 2e^2?

OpenStudy (stamp):

\[f'(x)=2e^{x+2}\]\[f'(0)=2e^{(0)+2}\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

how do you sqrt prime? let say f(x)=sqrt(4x)

OpenStudy (anonymous):

the final answer is 4x1/2?

OpenStudy (stamp):

u = 4x du/dx = 4 \[d(\sqrt {u})/du = \frac{1}{2\sqrt u}du/dx\]\[\frac{1}{2\sqrt {4x}}4=\frac{2}{\sqrt{4x}}=\frac{2}{2\sqrt x}=\frac{1}{\sqrt x}\]

OpenStudy (anonymous):

Ok,the final answer is 1/sqrt(x)